OpenStudy (samsan9):

use rules or logarithms to expand? log (x+3)^2/(x-2)(x^2+5)^4=

OpenStudy (samsan9):

\[\log \frac{ (x+3)^2 }{ (x-2)(x^2+5)^4 }\]

OpenStudy (samsan9):

idk what to do

OpenStudy (abmon98):

\[\log _{b}(m/n) = \log _{b}(m) – \log _{b}(n)\]

OpenStudy (samsan9):

what about the exponents?

OpenStudy (abmon98):

\[ \log _{b}(mn) = \log _{b}(m) + \log _{b}(n)\] \[2\log(x-3)-\log(x-2)+4\log(x^2+5)\]

OpenStudy (abmon98):

i am sorry i have to go now its 4:21 AM